A drawer contains exactly 5 red, 4 blue and 3 green pencils. If two pencils are selected at random, one after the other without replacing the first, what is the probability that the first one is red and the second one is green?
Answer by stanbon(75887)

A drawer contains exactly 5 red, 4 blue and 3 green pencils. If two pencils are selected at random, one after the other without replacing the first, what is the probability that the first one is red and the second one is green?
----
Ans: (5/12)(3/11) = 15/132 = 0.1136
====================================
Cheers,
Stan H.
==========